# Self-Tuning: Instructing LLMs to Effectively Acquire New Knowledge through Self-Teaching

### v4

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from unseen raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. Additionally, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on various models, e.g., Llama2-7B reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.
@@ 1 Introduction
Armed with a wealth of factual knowledge acquired during the pre-training phase sections, LLMs Touvron et al. (2023a); OpenAI (2023) exhibit remarkable proficiency in numerous knowledge-intensive tasks Cohen et al. (2023); Zhang et al. (2023); Gekhman et al. (2024). Despite this, the knowledge stored in LLMs can quickly become outdated due to the one-time training of LLMs and the ever-changing nature of the world Huang et al. (2023); Jiang et al. (2024c). These unavoidable knowledge limitations present notable obstacles to the trustworthiness of LLMs in real-world scenarios Liu et al. (2023); Mecklenburg et al. (2024). Thus, it is essential to equip LLMs with new knowledge to keep them up-to-date.
@@ 3.1 Document Collection and QA Pair Generation
To construct the document corpus, we collect articles from September to October 2023 (4,257 articles in total) from Wikipedia NewPages111https://en.wikipedia.org/wiki/Special:NewPages, which include new articles from various domains published after the pre-training cut-off time of the LLMs being evaluated.222The pre-training cut-off time for the Llama2 family models used in this study is 2022. Following Jiang et al. (2024c), we only use the first paragraph of each article, as it offers a comprehensive summary and contains a wealth of factual information.
@@ 4.2 Self-Teaching Learning Strategy
Gist identification improves the model’s ability to pinpoint the key elements (i.e., entities) within the atomic facts. Specifically, we prompt the model with Highlight the key information within the article:, and use the entities within the document as gold answers, identified using Spacy333https://spacy.io/usage.
Natural language inference provides the model with the capability to determine whether a statement can be inferred from specific document contents (i.e., “Yes,” “No,” or “It’s impossible to say”), thus avoiding misconceptions that may arise during knowledge acquisition. Specifically, we use a randomly sampled sentence (identified using NLTK444A natural language toolkit. https://www.nltk.org/) within the document content as the true statement, and a corrupted version where one entity is replaced by an irrelevant entity from another sentence as the false statement. Then, we prompt the model with Based on the article above can we conclude that and the sampled sentence (either initial or corrupted), with the three relations as options and corresponding answers.
@@ 6 Conclusion
Our study primarily focuses on enhancing a language model’s ability to effectively learn new knowledge from previously unseen raw corpora. Although experimental results on MCQA and NQ demonstrate that our Self-Tuning method well preserves previously acquired knowledge, future research could explore integrating Self-Tuning with continual learning approaches Wang et al. (2024). For instance, regularization-based methods such as EWC Kirkpatrick et al. (2017) and replay-based methods, like incorporating segments from general domain datasets (e.g., Wiki data Zhang et al. (2024c)), could improve the model’s capacity to retain learned knowledge and skills while mitigating the risk of overfitting to new information.
@@
Ethics Statement
Throughout our research, we have consistently adhered to ethical guidelines to uphold privacy, fairness, and the well-being of all individuals and groups involved. All benchmark datasets utilized in this study are used solely for research purposes and do not contain any personally identifiable information, thereby safeguarding privacy. During the QA data collection process for Wiki-Newpages-2023-QA, WikiBio-2024, and WebNews-2023 using GPT-4, we meticulously crafted prompts to eliminate any language that might discriminate against specific individuals or groups. These measures were implemented to minimize potential negative effects on users’ well-being. Examples of these carefully designed prompts can be found in Table 23, Table 24, and Table 26. To further ensure the quality of the newly collected datasets, the authors manually reviewed them following the guidelines in Bai et al. (2022). These datasets were confirmed to be of high quality, free from offensive content, false information, and any personally identifiable information Radharapu et al. (2023); Zhou et al. (2023). Future research efforts could explore the OpenAI moderation API555https://platform.openai.com/docs/guides/moderation/overview to systematically filter out inappropriate system responses. Additionally, GPT-4 is solely used to convert document knowledge into QA pairs—without introducing any additional information—and can be replaced by any other model.
@@ Appendix I Evaluation Results on Llama2-13B in the Single-domain Scenario
This section explores the potential of integrating Self-Tuning with continual learning techniques. Table 16 presents the evaluation results of combining Self-Tuning with a representative continual learning strategy: the replay-based method. In this approach, 500 training QA pairs were randomly sampled from the Wiki QA datasets, curated prior to the pre-training cutoff date of Llama2-7B Zhang et al. (2024c). These QA pairs were included throughout the training process to reinforce general domain knowledge.
@@
Future research could explore the inclusion of segments from general domain datasets, such as Wiki data Zhang et al. (2024c) and the Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021), which were compiled prior to the pre-training cut-off date. Adopting this strategy may improve the model’s capacity to retain learned knowledge and skills while reducing the risk of overfitting to novel information. In our current study, we deliberately avoid integrating extra data to ensure a precise assessment of knowledge injection, thereby preventing any biases that might arise from the inclusion of additional sources.

### v5

@@ Abstract
Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. To keep LLMs current, existing approaches typically involve continued pre-training on new documents. However, they frequently face difficulties in extracting stored knowledge. Motivated by the remarkable success of the Feynman Technique in efficient human learning, we introduce Self-Tuning, a learning framework aimed at improving an LLM’s ability to effectively acquire new knowledge from unseen raw documents through self-teaching. Specifically, we develop a Self-Teaching strategy that augments the documents with a set of knowledge-intensive tasks created in a self-supervised manner, focusing on three crucial aspects: memorization, comprehension, and self-reflection. Additionally, we introduce three Wiki-Newpages-2023-QA datasets to facilitate an in-depth analysis of an LLM’s knowledge acquisition ability concerning memorization, extraction, and reasoning. Extensive experimental results on various models, e.g., Llama2-7B reveal that Self-Tuning consistently exhibits superior performance across all knowledge acquisition tasks and excels in preserving previous knowledge.††∗Work done during the internship at Tencent AI Lab. 111Our source code and the curated new datasets can be accessed publicly at https://github.com/zhangxy-2019/Effective-Knowledge-Injection.
@@ 1 Introduction
Armed with a wealth of factual knowledge acquired during the pre-training phase Zhou et al. (2023); Zhang et al. (2025), LLMs Touvron et al. (2023a); OpenAI (2023) exhibit remarkable proficiency in numerous knowledge-intensive tasks Cohen et al. (2023); Zhang et al. (2024c, e); Gekhman et al. (2024). Despite this, the knowledge stored in LLMs can quickly become outdated due to the one-time training of LLMs and the ever-changing nature of the world Zhang et al. (2022, 2023); Huang et al. (2023); Jiang et al. (2024c). These unavoidable knowledge limitations present notable obstacles to the trustworthiness of LLMs in real-world scenarios Liu et al. (2023); Mecklenburg et al. (2024). Thus, it is essential to equip LLMs with new knowledge to keep them up-to-date.
@@ 3.1 Document Collection and QA Pair Generation
To construct the document corpus, we collect articles from September to October 2023 (4,257 articles in total) from Wikipedia NewPages222https://en.wikipedia.org/wiki/Special:NewPages, which include new articles from various domains published after the pre-training cut-off time of the LLMs being evaluated.333The pre-training cut-off time for the Llama2 family models used in this study is 2022. Following Jiang et al. (2024c), we only use the first paragraph of each article, as it offers a comprehensive summary and contains a wealth of factual information.
@@ 4.2 Self-Teaching Learning Strategy
Gist identification improves the model’s ability to pinpoint the key elements (i.e., entities) within the atomic facts. Specifically, we prompt the model with Highlight the key information within the article:, and use the entities within the document as gold answers, identified using Spacy444https://spacy.io/usage.
Natural language inference provides the model with the capability to determine whether a statement can be inferred from specific document contents (i.e., “Yes,” “No,” or “It’s impossible to say”), thus avoiding misconceptions that may arise during knowledge acquisition. Specifically, we use a randomly sampled sentence (identified using NLTK555A natural language toolkit. https://www.nltk.org/) within the document content as the true statement, and a corrupted version where one entity is replaced by an irrelevant entity from another sentence as the false statement. Then, we prompt the model with Based on the article above can we conclude that and the sampled sentence (either initial or corrupted), with the three relations as options and corresponding answers.
@@ 6 Conclusion
Our study primarily focuses on enhancing a language model’s ability to effectively learn new knowledge from previously unseen raw corpora. Although experimental results on MCQA and NQ demonstrate that our Self-Tuning method well preserves previously acquired knowledge, future research could explore integrating Self-Tuning with continual learning approaches Wang et al. (2024). For instance, regularization-based methods such as EWC Kirkpatrick et al. (2017) and replay-based methods, like incorporating segments from general domain datasets (e.g., Wiki data Zhang et al. (2024d)), could improve the model’s capacity to retain learned knowledge and skills while mitigating the risk of overfitting to new information.
@@
E Qualitative Analysis
F Ablation Study
G Error Analysis
H Evaluation Results on Self-Tuning Variants
I Evaluation Results on Llama2-13B in the Single-domain Scenario
J Evaluation Results on Llama2-7B-chat in the Single-domain Scenario
K Training Efficiency Analysis
L Evaluation Results on Gemma-7B in the Single-Domain Scenario
M Evaluation Results with Continual Learning Techniques
N Evaluation Results Comparing with a Baseline Utilizing Document-Based QA Generation on Test Corpora
- O In-depth Sample Documents and Corresponding QA Pairs for Open-Ended Generation and Natural Language Inference Tasks
N Evaluation Results Comparing with a Baseline Utilizing Document-Based QA Generation on Test Corpora
M Evaluation Results with Continual Learning Techniques
L Evaluation Results on Gemma-7B in the Single-Domain Scenario
K Training Efficiency Analysis
J Evaluation Results on Llama2-7B-chat in the Single-domain Scenario
I Evaluation Results on Llama2-13B in the Single-domain Scenario
H Evaluation Results on Self-Tuning Variants
G Error Analysis
F Ablation Study
6 Conclusion
5.6 Results of Varying Models and Corpora
@@ Appendix I Evaluation Results on Llama2-13B in the Single-domain Scenario
This section explores the potential of integrating Self-Tuning with continual learning techniques. Table 16 presents the evaluation results of combining Self-Tuning with a representative continual learning strategy: the replay-based method. In this approach, 500 training QA pairs were randomly sampled from the Wiki QA datasets, curated prior to the pre-training cutoff date of Llama2-7B Zhang et al. (2024d). These QA pairs were included throughout the training process to reinforce general domain knowledge.
@@
Future research could explore the inclusion of segments from general domain datasets, such as Wiki data Zhang et al. (2024d) and the Massive Multitask Language Understanding (MMLU) Hendrycks et al. (2021), which were compiled prior to the pre-training cut-off date. Adopting this strategy may improve the model’s capacity to retain learned knowledge and skills while reducing the risk of overfitting to novel information. In our current study, we deliberately avoid integrating extra data to ensure a precise assessment of knowledge injection, thereby preventing any biases that might arise from the inclusion of additional sources.